# Emergence of global receptive fields capturing multipartite quantum correlations

### v1

@@ I.1 Permutation structure and correlations of the Dicke states
We start by analyzing the properties of the real-valued non-negative wave functions that belong to the Dicke state family,
@@
Figure 2: Ursell functions up to 4th order that show the correlations structure of the 16-qubit Dicke states. The labels such as $xxzz$ denote spin projection indexes of the calculated correlation functions. In the case of the 3rd and 4th order Ursell functions the correlation levels are multiply degenerate and for presentation purposes each level is associated only with one of correlation functions. The complete correlation histograms are presented in Appendix C. $all$ stands for the cases when all the correlation functions are zero.
The difference in structure of the permutations used for constructing |$\Ket{\Psi^{D}_{N}}$ with different $D$ parameters is intimately related to variance of quantum correlations of these functions. The previous consideration Multipartite_correlations of the quantum correlations on the level of the density matrices with relative entropy has demonstrated that the Dicke states with $D=\frac{N}{2}$ display correlations at any order of a coarse grained partition of the system in question, which evidences a complex multipartite structure of these wave functions. From the perspective of applying a quantum state neural network approach in which one is to approximate unknown probability distribution for the basis functions of a target state, it is instructive to analyze the classical averages that are the correlation functions between qubits. For that we calculate the Ursell functions Ursell (connected correlations) up to fourth order for spin operators, ${\hat{\sigma}}_{i}^{\alpha}$, where $i$ is the site index and $\alpha$ stands for the projection $x$, $y$ or $z$. In our case we concentrate on the first order local correlations, $\Gamma^{\alpha}_{i}$ and non-local correlation functions of higher orders, $\Gamma^{\alpha\beta}_{ij}$, $\Gamma^{\alpha\beta\gamma}_{ijk}$ and $\Gamma^{\alpha\beta\gamma\delta}_{ijkl}$ ($i\neq j\neq k\neq l$). The corresponding expressions for these correlation functions are given in Appendix C (Fig. 7).
@@
I.2 Learning correlations with global receptive fields
We are now in position to discuss the neural quantum state tomography for the Dicke states, Eq.1 and to show that the receptive fields capturing non-local correlations can be discovered within the training without imposing symmetries. In the tomography approach the target wave function is to be approximated with positive neural quantum state
where $N$ ($M$) is the number of the visible (hidden) units, $\boldsymbol{\lambda}=\{{\bf a},{\bf b},{\bf W}\}$ denotes the parameters of the restricted Boltzmann machine and the partition function $Z_{\boldsymbol{\lambda}}$ ensures the normalization of $p_{\boldsymbol{\lambda}}$. $W_{ij}$ is the weight matrix between visible and hidden units and $a_{i}$ ($b_{j}$) is the bias of the $i$th visible ($j$th hidden) neuron. The goal of the tomographic procedure is to optimize the network weights in such a way that the approximated wave function, Eq.3 has the largest overlap with the target wave function that is one of the Dicke states, Eq.1. For these purposes, one uses the Kullback-Leibler divergence Amari as the loss function. Importantly, the information about the target wave function is only represented in the form of finite sets of the bitstrings obtained from projective measurements. The comprehensive description of the neural quantum state tomography procedure we use can be found in Refs. qucumber ; Torlai_thesis ; NNtomography .
Previously, the real-valued Dicke states with $D=1$ (known as the $W$ state) for the systems of 20, 40 and 80 qubits were reconstructed with the neural quantum state tomography NNtomography . Since the |$\Ket{\Psi^{1}_{N}}$ is the simplest wave function in the Dicke family and is characterized by the smallest entanglement, the RBM used for the tomography contains the same number of visible and hidden units. The reported overlap between the ideal $W$ state and its RBM approximation for $N=20$ qubits is very high, 0.997. To our knowledge, the tomographic results for the other states with larger $D$ that belong to the family have not been demonstrated. At the same time by increasing the Dicke index $D$ from 1 to $\frac{N}{2}$, one increases the complexity of the quantum state as discussed in the previous section. In our work we focus on these highly-entangled states.
Fig. 3 visualizes the neural quantum state tomography results which are the elements of the weight matrices, $W_{ij}$ between visible and hidden neurons of the RBMs we used to approximate different Dicke states for 8-, 16 and 20-qubit systems. The resulting weight matrices are square, since in all the cases the number of hidden and visible neurons is the same. One can see that for the simplest wave functions with $D=1$ the hidden units show random-like patterns of interactions with neurons from the visible layer. In this case absolute values of the $W$ matrix elements do not exceed 9. In agreement with the results of Ref. NNtomography, the target states with $D=1$ are reconstructed with high fidelity of 0.99 independent on the system size.
Remarkably, the solutions of the tomographic problem obtained for more entangled wave functions with $D>1$ reveal distinct patterns of the elements of the weight matrices. Each hidden unit is characterized by the receptive fields having non-zero weights with all the input neurons. There is a maximal (positive) or minimal (negative) interaction with the specific visible neuron, which indicates weight sharing and is an important feature of convolution CNN-LeCun . Weak non-zero couplings to all other neurons from visible layer (Fig.3 D) reflect a distinct type of correlations of the input data. Thus, in the case of the RBMs approximating the Dicke states with large $D$, the whole weight space is uniformly covered with such global RFs.
The results presented in this section were obtained for 8-, 16- and 20-qubit systems by using the optimization procedure as described in Ref.NNtomography with the same number of input and hidden neurons. As one would expect, the increase of the Dicke index $D$ leads to a considerable degradation of reconstructing the quantum states (Fig.3), and the number of the hidden units should be substantially increased so that the overlap is close to 90% (Appendix A). Below we show that when constructing RBM to approximate a Dicke wave function one can work around the hard problem of the neural quantum tomography by utilizing the architecture constructed with the revealed global receptive fields.
I.3 A compact and exact RBM representation for all $N$-qubit Dicke states with global RFs
Searching for exact classical neural network representations of complex entangled quantum states is one of the challenging tasks in modern science Duan_representation ; Imada_representation ; Ivan_Glasser . If possible, this avoids performing expensive training procedure and saves computer resources by compressing information about a wave function using the weights of a neural network. In this sense, the Dicke states are no exception. As shown in Ref.NNtomography the |$\Ket{\Psi^{D}_{N}}$ wave function can be represented exactly with RBM, however, for that one should reserve single hidden unit for each pair of qubits in the system in question. This means the total number of parameters grows as $N^{2}$ as the system size increases. The most compact representation of a $N$-qubit Dicke state with arbitrary $D$ in terms of the number of hidden neurons so far has been reported in Ref.Clark1 . In this work the author has used a correlator product state approach to encode |$\Ket{\Psi^{D}_{N}}$ and analytically showed that any Dicke wave function can be described with $M=[\frac{N}{2}]$. Here, we are going to demonstrate that by using global receptive fields one can develop exact and compact RBM representation for the whole Dicke wave functions family with $M=N$.
To imitate the Dicke states with RBM, we employ a non-sparse circular square weight matrix, $W$ that is schematically presented in Fig.4 and is filled with only two real-valued parameters, $W_{\rm min}$ and $W_{\rm max}$. Taking into account the structure of the discovered global receptive field, in our model RBM each hidden neuron has the $W_{\rm max}$ interaction with one unique visible unit and the $W_{\rm min}$ with other visible sites. Importantly, these weights are of different signs, $W_{\rm min}<0$ and $W_{\rm max}>0$. Since $M=N$, $W$ is the square matrix and according to Ref.Clark2 our RBM can be classified as compact one. The biases for visible and hidden neurons are set to zero, $\mathbf{a}=\mathbf{b}=0$. Constraining weight matrix structure in such a way makes the amplitudes of the basis states with the same number of ”1”s to be the same. In other words, all the basis functions can be distinguished by the parameter $d=\sum_{i}v_{i}$. For the given $W_{\rm min}$ and $W_{\rm max}$ the basis functions, |$\Ket{\bf v}$ that are characterized by the same $d$ have the same amplitude
where $d[{\bf v}]=\sum_{i}v_{i}$.
To construct the $W_{\rm min}$ $W_{\rm max}$ phase diagram that distinguishes the RBM states with respect to the Dicke wave functions with different indices $D$, the continuous parameter space is first discretized with a mesh. Then, for each point of the mesh the fidelities, $\mathcal{F}_{D}$ between the neural quantum state |$\Ket{\Psi_{\boldsymbol{\lambda}}}$ and all the Dicke states are calculated. More specifically, one defines a set of fidelities $\mathcal{F}_{D}=|\braket{\Psi^{D}_{N}}{\Psi_{\boldsymbol{\lambda}}}|^{2}$, where $0\leq D\leq N$. In general case, such fidelity calculations require determination of the normalization constant, $Z_{\boldsymbol{\lambda}}$ and summation over the whole basis states, which cannot be performed for large number of qubits $N$. Fortunately, the proposed architecture of the neural quantum state allows to directly estimate its fidelity with any Dicke state by means of the following expression:
It is also possible to derive a condition on $W_{\max}>0$ and $W_{\min}<0$ that maximizes the fidelity for a given Dicke state. Consider the case, in which $K$ of $N$ visible neurons are active, and let us find the optimal configuration of hidden neurons that maximizes the system’s energy. Enabling any hidden neuron that is not connected to $K$ active visible neurons decreases the total energy by $KW_{\min}$, so these neurons are preferably inactive. At the same time, enabling $K$ remaining hidden neurons causes a shift in energy of $\Delta E=W_{\max}+(K-1)W_{\min}$. Given $\Delta E>0$, the energy will be maximized by activating all $K$ hidden neurons connected to $K$ activated visible neurons. The total energy in this case is $K\Delta E=KW_{\max}+K(K-1)W_{\min}$. The maximum over $K$ is obtained when $K=(1-W_{\max}/W_{\min})/2$. In this way, unit fidelity for |$\ket{\Psi_{N}^{D}}$ is achieved when
$\frac{W_{\max}}{-W_{\min}}=2D-1$ (7)
II Discussions and perspectives
Deep convolution networks with local receptive fields are known to be computationally inefficient for capturing large-scale correlations in data of different origin, which stimulates developing distinct types of the neural models such as non-local neural networks Wang and transformers transformer_original . Constructing non-local receptive fields to extend the capabilities of CNN and other architectures is problem specific and hard to do from scratch. However, our study devoted to approximation of the entangled Dicke wave functions shows that one can employ additional unsupervised learning tools such as quantum neural state tomography and, in this way, get basic information about optimal structure of the receptive fields within the learning procedure. Remarkably, based on our results concerning the Dicke states such a learning does not have to be perfect.
As an interesting direction for future investigations we consider searching for receptive fields when performing neural quantum state tomography on the basis of simple and deep RBMs for wave functions characterized by different amount and structure of entanglement, for instance Greenberger–Horne–Zeilinger state, the tensor product of maximally entangled two-qubit states and others. In the same spirit, one can analyze the weight spaces of the trained RBM approximating ground states of quantum Hamiltonians on regular lattices, for instance one dimensional Ising model and Heisenberg Hamiltonian on the square lattice and others. In some cases neural network models of the aforementioned wave functions are known Clark1 ; Clark2 ; NNtomography ; Duan_representation ; Imada_representation ; Ivan_Glasser and we expect to find receptive fields of different shapes and properties, which facilitates developing distinct compact classical representations for these quantum states through learning.
Acknowledgements
Appendix A Neural quantum state tomography of the Dicke states
In this section we describe the details of the neural quantum state tomography whose results are presented in the main text. All the simulations presented in the main text were performed with the following parameters. We used 10 steps within the contrastive divergence procedure and the learning rate was taken to be 0.1. The maximal number of the training epochs was set to 2000. However, in some calculations we observed the degradation of the results concerning fidelity with the target quantum state. That is why the neural network weights corresponding to the best fidelity have been selected for further analysis.
Figure 6: Fidelity between ideal |$\Ket{\Psi^{8}_{16}}$ Dicke wave function and its RBM approximations obtained within neural quantum state tomography for different number of the hidden neurons. In these numerical experiments the learning rate was taken to be 0.01.
The results presented in Fig.3 in the main text were obtained with the same number of the hidden and visible neurons. Since the complexity of the quantum states enhances as the Dicke index increases we observe the degradation in reconstructing the target state as described in the main text. This problem can be solved by increasing the number of the hidden neurons and decreasing the learning rate. From Fig.6 plotted for the |$\Ket{\Psi^{8}_{16}}$ state characterized by the largest entanglement it follows that the quality of the neural network approximation can be slightly improved by adding more hidden neurons, which likewise confirms the importance of the RF approach we develop in this work.
Appendix B Derivation of the fidelity expression
where $p_{\boldsymbol{\lambda}}(d)$ are normalized RBM probabilities:
Then, one can write the following expression for fidelity:
Substitution of (9) in (10) gives equation (6) in the main text.
Appendix C Ursell functions

### v2

@@ Appendix A Permutation structure and correlations of the Dicke states
In this appendix we analyze the properties of the real-valued non-negative wave functions that belong to the Dicke state family,
@@
The difference in structure of the permutations used for constructing |$\Ket{\Psi^{D}_{N}}$ with different $D$ parameters is intimately related to variance of quantum correlations of these functions. The previous consideration Multipartite_correlations of the quantum correlations on the level of the density matrices with relative entropy has demonstrated that the Dicke states with $D=\frac{N}{2}$ display correlations at any order of a coarse grained partition of the system in question, which evidences a complex multipartite structure of these wave functions. From the perspective of applying a quantum state neural network approach in which one is to approximate unknown probability distribution for the basis functions of a target state, it is instructive to analyze the classical averages that are the correlation functions between qubits. For that we calculate the Ursell functions Ursell (connected correlations) up to fourth order for spin operators, ${\hat{\sigma}}_{i}^{\alpha}$, where $i$ is the site index and $\alpha$ stands for the projection $x$, $y$ or $z$. In our case we concentrate on the first order local correlations, $\Gamma^{\alpha}_{i}$ and non-local correlation functions of higher orders, $\Gamma^{\alpha\beta}_{ij}$, $\Gamma^{\alpha\beta\gamma}_{ijk}$ and $\Gamma^{\alpha\beta\gamma\delta}_{ijkl}$ ($i\neq j\neq k\neq l$). The corresponding expressions for these correlation functions are given in Appendix B.
@@
Appendix B Ursell functions
@@
Appendix C Neural quantum state tomography of the Dicke states
In this section we describe the details of the neural quantum state tomography whose results are presented in the main text. All the simulations presented in the main text were performed with the following parameters. We used 10 steps within the contrastive divergence procedure and the learning rate was varied in the range [0.01, 0.2] to reach the best possible fidelity with target state. The number of the training epochs was set to 10000. In some calculations we observed the degradation of the results concerning fidelity with the target quantum state with number of epochs. That is why the neural network weights corresponding to the best fidelity within a run have been selected for further analysis.
Fig.9 demonstrates the quality of RBM approximation that has been achieved with different number of the bitstrings in the training set, $N_{s}$ in the case of the 16-qubit Dicke wave functions. In agreement with Ref.NNtomography for $D=1$ ($W$ state) the fidelity approaches to 1 at $10^{3}$ samples, which is due to a strong localization of the target state in the Hilbert space. In the case of the $D=8$ state that is characterized by the largest entanglement for 16-qubit system the saturation value of the fidelity with the exact solution is about 0.7. As follows from Fig.9, increasing the number of bitstrings in the training set does not improve the quality of the tomographic procedure.
Figure 9: Fidelity between the exact 16-qubit Dicke states and their approximations generated by trained RBMs. Grey squares and blue circles denote results obtained for the cases of $D=1$ and $D=8$, respectively. The number of the hidden units was fixed in these simulations as $M=N=16$.
Figure 10: Fidelity between ideal 16-qubit Dicke wave functions and their RBM approximations obtained within neural quantum state tomography for different number of the hidden neurons.
Figure 11: Fidelity between ideal |$\Ket{\Psi^{8}_{16}}$ Dicke wave function and its RBM approximations obtained within neural quantum state tomography for different number of the hidden neurons.
The results presented in Fig.9 and Fig. 3 in the main text were obtained with the same number of the hidden and visible neurons. To complete this consideration it is important to explore the regimes when $M<N$ and $M>N$. 16-qubit examples of the former are given by Fig.10. One can see that for $D=1$ the fidelity of 0.99 with exact solution can be reached with 10 hidden neurons. In turn, quality of the |$\ket{\Psi^{8}_{16}}$ approximation is rather low and the corresponding fidelity takes the values of about 0.7 for $M>14$. As described in the main text such a degradation of the fidelity when varying the Dicke index is due to the complexity of the target quantum states that enhances as $D$ increases. This problem can be solved by increasing the number of the hidden neurons (the $M>N$ regime). From Fig. 11 plotted for the |$\Ket{\Psi^{8}_{16}}$ state characterized by the largest entanglement it follows that the quality of the neural network approximation can be slightly improved by adding more hidden neurons, which likewise confirms the importance of the RF approach we develop in this work.
To conclude this section by the example of Fig.12 we provide an evidence that RF field can emerge even when one employs a RBM with one or a few hidden neurons. Despite of a low quality of the resulting approximations (in both cases the overlap with the exact state is about 0.2) one can clearly distinguish the profile of the global receptive field formed in the parameter space of the trained RBM with minimal $M$. This leads us to an important conclusion that an optimal network structure could be also developed on the basis of preliminary fast learning with a minimal network setting.
Figure 12: Formation of the global receptive fields in the RBMs with a few hidden neurons. (top) and (bottom) RBM weights between hidden and visible neurons obtained within the neural quantum state tomography of the |$\ket{\Psi^{8}_{16}}$. The RBMs feature one (top) and three (bottom) hidden units. The fidelity between RBM approximation and target wave function is 0.20 (top) and 0.22 (bottom).
Appendix D Derivation of the fidelity expression
where $p_{\boldsymbol{\lambda}}(d)$ are normalized RBM probabilities:
Then, one can write the following expression for fidelity:
Substitution of (23) in (24) gives equation (6) in the main text.